OpenStudy (riley):

Convert the exponential function to the form indicated. (Round all coefficients to four significant digits.) f(x) = 2.5 e^(-0.7 x); f(x)=A b^x A = ? b = ?

OpenStudy (anonymous):

f(x) = 2.5 * e^(-0.7x) f(x) = 2.5 * (e^-0.7)^x A=2.5 b=e^-0.7=0.4965

OpenStudy (riley):

how did you do that o_O?

OpenStudy (anonymous):

Are you referring to this type of transformation, or something else I did ?